(24/(√3*(5-2)))/[-1-(-3)]^2
Answer by CubeyThePenguin(3113)

(24/(√3*(5-2)))/[-1-(-3)]^2
= (24/(sqrt(3 * 3))/((2)^2)
= (24/3)/(4)
= 8/4
= 2